jim has scores of 60,73,82, & 75.
He wants to raise the mean to 75.

Answer by unlockmath(1688)

Hello,
With this we can set it up as:
(60+73+82+75+x)/5 = 75
Multiply by 5:
60+73+82+75+x = 375
add like terms:
290+x=375
Subtract 290:
x=85
Make sense?
RJ
www.math-unlock.com